OpenStudy (anonymous):

int_{0}^{2\pi} cos^2(x) dx

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi} \cos^2(x) dx \]

OpenStudy (anonymous):

Use the identity \(\cos^2x=\dfrac{1+\cos2x}{2}\)

OpenStudy (anonymous):

got it thanks.

OpenStudy (anonymous):

Got pi